# Towards Pervasive Distributed Agentic Generative AI - A State of The Art

### v1

@@ Abstract.
The rapid advancement of intelligent agents and Large Language Models (LLMs) is reshaping the pervasive computing field. Their ability to perceive, reason, and act through natural language understanding enables autonomous problem-solving in complex pervasive environments, including the management of heterogeneous sensors, devices, and data. This survey outlines the architectural components of LLM agents (profiling, memory, planning, and action) and examines their deployment and evaluation across various scenarios. Than it reviews computational and infrastructural advancements (cloud to edge) in pervasive computing and how AI is moving in this field. It highlights state-of-the-art agent deployment strategies and applications, including local and distributed execution on resource-constrained devices. This survey identifies key challenges of these agents in pervasive computing such as architectural, energetic and privacy limitations. It finally proposes what we called ”Agent as a Tool”, a conceptual framework for pervasive agentic AI, emphasizing context awareness, modularity, security, efficiency and effectiveness.
@@
ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language generationccs: Human-centered computing Ubiquitous computingccs: Networks Cloud computing
@@ 1. Introduction
Agents, designed to perceive, reason, and act within their environment, are pivotal for solving complex tasks. The advent of Large Language Models (LLMs) has recently revolutionised this domain. Trained on vast web datasets, LLMs demonstrate an impressive understanding of human language and can generate remarkably similar and accurate responses. Integrating these capabilities has led to the development of LLM-based agents, where LLMs serve as the core cognitive engine, combined with perceptual, reasoning, and action mechanisms. This synergy has introduced a new paradigm of more intelligent and versatile agents applicable across diverse domains. Agents can have various capabilities. Some are purely reactive, responding to explicit user prompts (Franciscatto et al., 2025; Zhang et al., 2024e), while others show proactive functionalities by autonomously initiating tasks based on their understanding of the environment (Zhang et al., 2024i). Their applications are widespread. Embodied agents, interact with the physical world via sensors and actuators (Durante et al., 2024), whereas software agents operate in digital environments, performing tasks like information retrieval or software and research development (Zhang et al., 2024h, d).
Also the field of pervasive computing, focused on integrating computing and communication environments into human daily life, has made considerable progress. This ranges from high-performance computing (HPC) systems to resource-constrained IoT devices, with applications in smart homes, cities, factories, or hospitals. In this pervasive context, agents offer vast potentials. Their abilities to perceive, reason, and interact can significantly enhance the scope and utility of pervasive computing applications. However, deploying agents in these environments presents significant computational, and security challenge that necessitate custom solutions. Researchers have proposed various solutions to solve these challenges. Therefore, this paper analyses strategies for integrating agents into pervasive computing, examines existing applications, identifies persistent challenges, and explores ongoing research efforts to address them. Specifically, the paper is structured as follows:
@@
    First, an introduction to LLM-based agent systems will cover their architecture, performance evaluation metrics for accuracy and robustness, and diverse applications.
@@ 2. Agent LLM Architecture
As described by Franklin and Graesser, an agent can be defined as a system that exists within an environment, it perceives that environment through observation using sensors, and acts upon it through effectors, over time, in pursuit of its own agenda and so as to effect what it senses in the future (Franklin and Graesser, 1996).
In the context of LLM-based agents, a natural language processing pipeline is employed to enable the agent to perceive, reason, and interact with its environment to accomplish specific goals. Perception, the initial stage in this process, involves gathering information about the current state of the environment (Wang et al., 2024c). This often occurs through natural language inputs, where users provide instructions or queries (Franciscatto et al., 2025), or structured text, including HTML or a programming language formats (Hu et al., 2024; Zhang et al., 2024e). In addition, multimodal LLMs (MLLM), process visual and auditory data, resulting in enhanced environmental comprehension and information use (Durante et al., 2024). Also, to enhance environmental interaction, respect constraints, and achieve goals, LLM-based agents require a robust understanding of implied textual meanings (Li et al., 2024a). Following perception, the agent leverages the LLM’s natural language processing capabilities for reasoning and planning. This involves in processing the perceived information, interpreting user intent, and analyzing possible courses of action (Parmar et al., 2025; Wang et al., 2024c). A key aspect of this stage is the formulation of strategic plans to achieve desired outcomes. Complex tasks are often decomposed into manageable subtasks to facilitate execution (Li et al., 2024a; Parmar et al., 2025) such as for robots movements management (Yang et al., 2025a), automatic world exploration (Wang et al., 2023) or code debugging (Wang et al., 2024b). The final stage involves in interacting directly with the environment following the previous reasoning steps (Wang et al., 2024c). Actions may include generating natural language responses to the user, providing explanations or completing tasks through dialogue (Zhang et al., 2024d; Xi et al., 2025), invoke external tools or APIs (Durante et al., 2024; Schick et al., 2023) or produce executable code (Zhang et al., 2024h). To accomplish all these tasks modern agent architectures typically consist of multiple key modules that work together to create intelligent behavior. A typical LLM-agent architecture has four key modules: a profiling module for defining the agent’s role, a memory module for storing and retrieving past experiences, a planning module for formulating future actions, and an action module for translating decisions into outputs (Zhang et al., 2024d). Also, evaluating this architecture is essential to determine the effectiveness of its modules, especially within the intended application domain. The following sections will explore the LLM architecture, the specific functionalities and interactions of the agent modules, and the evaluation strategies and application domains of these intelligent LLM agents.
@@ 2.1. Large Language Models
Figure 1. The transformer architecture, from (Vaswani et al., 2017)
The LLM architecture serves as the core cognitive component, or ”brain”, of these intelligent agents. Its fundamental operation relies on ”next token prediction”. This means that given an input sequence of words (tokens), transformed into numerical embeddings, the LLM predicts the subsequent word in the sequence, iteratively generating the output. The success of LLMs in this task is largely attributed to their foundation in the Transformer architecture, illustrated in Figure 1. Its design captures long-range dependencies within the text and effectively models the complex structures of human language (Vaswani et al., 2017). A typical LLM consists of stacked Transformer layers, each incorporating a multi-head self-attention mechanism and a position-wise fully connected feed-forward network. The multi-head self-attention mechanism allows the model to weigh the importance of different parts of the input sequence when processing each word. It processes the encoded input as a set of key-value pairs (K, V), where both keys and values have a dimension equal to the input sequence length (n). The Transformer employs scaled dot-product attention to calculate how much attention each value should receive based on a query (Q):
@@
LLMs commonly employ this Transformer architecture, scaling it to billions or even trillions of parameters and incorporating various optimization strategies. For instance, decoder-only architectures like GPT-3 and LLaMA predict each next token based on the preceding ones. Conversely, encoder-only architectures such as BERT and RoBERTa prioritise understanding the input text to generate task-specific outputs like labels or token predictions (Matarazzo and Torlone, 2025). In the agentic field, the choice of architecture depends on the required task. Typically, decoder-only architectures are most commonly used when we want an agent to perceive, reason, and interact with the environment. Furthermore, Multimodal Language Models (MLLMs) are increasingly gaining traction for agents interacting in multimodal environments (text, images, video, audio). This is because they allow the application of Transformer architectures across diverse modalities like vision and audio. To enable such multimodal capabilities, MLLMs extend the Transformer architecture by embedding non-text inputs (e.g., image patches or audio tokens) into the same vector space as textual tokens. These embeddings are then processed jointly within shared Transformer layers or through modality-specific experts, such as in Multiway Transformer architectures. Deep modality fusion is achieved via cross-attention or unified encoders, enabling fine-grained interactions between visual and textual features (Liang et al., 2024).
@@ 2.2. Agent Modules
LLM-based autonomous agents are designed to effectively perform diverse tasks by leveraging the capabilities of Large Language Models (LLMs). To achieve this, their architecture typically incorporates 4 key modules: profiling module, memory module, planning module, and action module.
@@ 2.2.1. Profiling
Figure 2. An example of profiling prompt used in (Lu et al., 2024)
The primary purpose of the Profiling Module is to define the role-specific identity of the agent, which significantly influences its behavior and interactions (Wang et al., 2024c; Zhang et al., 2024d). This module is crucial for enabling agents to generate more contextually relevant responses from the LLM across different modalities. These profiles are typically added into the model context window. By including role descriptions in the prompt, the profiling module guides the LLM’s reasoning style, response formulation, and interaction strategy (Zhang et al., 2024d). For instance, in (Lu et al., 2024) the agent is profiled to work as PHD student that want to publish some papers (Figure 2). The profiling module serves as the foundation for agent design, giving a significant influence on the entire agent’s modules ecosystem (Wang et al., 2024c). The chosen profile can impact how the agent remembers information (memory module), how it formulates strategies to achieve its goals (planning module) that consequently will impact also the future actions. For example, an agent profiled as ’an expert machine learning researcher’ (Hu et al., 2024) might prioritise certain experiments over others or selectively memorise results more significant to its research. Similarly, an agent profiled as ’an advanced AI system serving as an impartial judge for intelligent code generation outputs’ (Zhuge et al., 2024) would focus more intently on code details, errors, and bugs than on other capabilities. Agent profiles are commonly implemented by handcrafting specific prompts that describe the desired role and behavior (Wang et al., 2024c). However, other strategies can be used, including LLM-generation of profiles and dataset alignment to reflect real-world characteristics (Wang et al., 2024c). Combining these strategies can yield additional benefits. Moreover agent profiles can include multiple dimensions, including basic attributes, behavioral patterns, and social information (Zhang et al., 2024d).
@@ 2.2.2. Memory
Figure 3. From left to right, the diagram displays the structures of Short-term, Long-term, and Hybrid memory.
@@
LLM architectures lack persistent memory across test time interactions, requiring prior exchanges to be explicitly included within the context window (Figure 3). This limitation underscores the need for effective memory management strategies in LLM-based agents, particularly for prioritizing relevant information during storage and retrieval. As the agent’s interaction history expands, exceeding the fixed context window’s capacity, summarization or information filtering becomes unavoidable. Consequently, autonomous LLM-based agents rely on a memory module to retain, recall, and reflect upon past observations, thoughts, feedback, and actions. This enables them to better understand the current context, anticipate future needs, correct their action trajectory, adjust their behavior, and maintain consistent and coherent behavior over time (Wen et al., 2024a; Xi et al., 2025; Wang et al., 2024c). This module often incorporate different structures to manage information across varying timescales:
@@
  • Short-Term Memory: This type of memory maintains contextually relevant information about recent perceptions, reasoning and actions (the trajectory history) within the LLM context window (Zhang et al., 2024d; Shinn et al., 2023). It allows for real-time adaptation during an interactive session. Examples include the internal states maintained by a conversational agent or the intermediate steps during a web research (Zhang et al., 2024e). However, the limited context window of LLMs poses a challenge for relying solely on short-term memory.
@@
  • Long-Term Memory: In contrast to the real-time encounter observations, which provide short-term memory, the overall encounter history represents long-term memory (Zhang et al., 2024a). Long-term memory can provide stable knowledge to complement the flexibility of short-term memory. This module is important because stores both successful experiences, demonstrating correct goal achievement strategies, and unsuccessful experiences, highlighting incorrect paths to avoid. For instance in (Wu et al., 2024a), successful paths are stored as ’Thought Cards’ and retrieved upon encountering a new question with a similar topic or in (Zhang et al., 2024i) the agent store historical events that will be used to predict potential tasks.
@@
  • Hybrid Memory: Many agents use a combination of short-term and long-term memory to leverage the benefits of both (Wang et al., 2024c). Short-term memory handles immediate context, while long-term memory provides access to a broader history and consolidated knowledge.
@@
These structures can be stored in various formats, each with its own advantages. One approach is to keep everything in natural language form. This is especially useful for conversational agents, where it’s important to be transparent about the agent’s history. For instance, in (Shinn et al., 2023) store past trials as verbal descriptions, making it easy to see what the agent has learned. Similarly, in (Wang et al., 2023) descriptions of skills within the Minecraft game are stored directly as natural language within the agent’s memory. Then, there’s the idea of Embedding Memory. Here, past experiences are transformed into high-dimensional vectors (Wang et al., 2024c). This allows for really fast reflection and helps the agent generalise across different situations. It’s especially good for searching through knowledge quickly, based on semantic similarity (Zhang et al., 2024d). Finally databases can be used to store memories. This gives us considerable power to manipulate the data through structured queries. So agents can be used to understand and execute SQL queries in natural language, which lets it interact with the database really effectively (Shen et al., 2024). Memory module supports three core operations. First writing, which involves persisting relevant information from past interactions into the memory. Key considerations during memory writing include handling duplicated information and managing memory overflow (Chen et al., 2024b). Techniques such as summarizing similar information, using a FIFO buffer to overwrite old entries (Wang et al., 2024c), truncating historical data that exceeds processing capacity (Xi et al., 2025), and implementing time limits to discard redundant data and prevent excessive memory consumption (Zhong et al., 2024) are used. Then, memory reading is a crucial operation that aims to obtain relevant knowledge from the agent’s memory to adapt its behaviors and inform its next actions (Zhang et al., 2024d). The retrieval of valuable information often considers factors like recency, relevance, and importance of the memories. These factors commonly include the recency of the memory, its relevance to the current situation or query, and its perceived importance (Xi et al., 2025; Wang et al., 2024c). Finally, memory reflection allows the agent to analyse, refine, and optimise the accumulated memories (Zhang et al., 2024d). It emulates human cognitive processes of summarizing and inferring more abstract and high-level insights from past experiences. Reflection can occur hierarchically, generating insights based on existing insights (Shinn et al., 2023). As mentioned in the previous section, the memory module is not isolated. It is influenced by the agent’s profile, which can determine what types of information are prioritised for storage and retrieval. Furthermore, the memory module plays a critical role in the planning process, providing the necessary context and historical data for the agent to formulate effective future actions. The planning module might retrieve information from memory to inform its reasoning and decision-making.

### v2

@@ Abstract.
The rapid advancement of intelligent agents and Large Language Models (LLMs) is reshaping the pervasive computing field. Their ability to perceive, reason, and act through natural language understanding enables autonomous problem-solving in complex pervasive environments, including the management of heterogeneous sensors, devices, and data. This survey outlines the architectural components of LLM agents (profiling, memory, planning, and action) and examines their deployment and evaluation across various scenarios. Then it reviews computational advancements (cloud to edge) in pervasive computing and how AI is moving in this field. It highlights state-of-the-art agent deployment strategies and applications, including local and distributed execution on resource-constrained devices. This survey identifies key challenges of these agents in pervasive computing such as architectural, energetic and privacy limitations. It finally proposes what we called ”Agent as a Tool”, a conceptual framework for pervasive agentic AI, emphasizing context awareness, modularity, security, efficiency and effectiveness.
@@
$c$s: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language generationccs: Human-centered computing Ubiquitous computingccs: Networks Cloud computing
@@ 1. Introduction
Agentic generative AI systems integrate generative foundation models (e.g., large language and multi-modal models) as the core reasoning engines inside autonomous agents. Designed to perceive, reason, and act within dynamic environments, these agents are pivotal for the successful decomposition and execution of complex, high-level objectives in areas previously inaccessible to static models. The advent of Large Language Models (LLMs) has recently revolutionised this domain. Trained on vast web datasets, LLMs demonstrate an impressive understanding of human language and can generate remarkably human-like and accurate responses. Integrating these capabilities has led to the development of LLM-based agents, where LLMs serve as the core cognitive engine, combined with perceptual, reasoning, and action mechanisms. This synergy has introduced a new paradigm of more intelligent and versatile agents applicable across diverse domains. Agents can have various capabilities. Some are purely reactive, responding to explicit user prompts (franciscatto2025cbr; zhang2024webpilotversatileautonomousmultiagent), while others show proactive functionalities by autonomously initiating tasks based on their understanding of the environment (zhang2024proagent). Their applications are widespread. Embodied agents interact with the physical world via sensors and actuators (durante2024agent), whereas software agents operate in digital environments, performing tasks like information retrieval or software and research development (zhang2024aflow; zhang2024agentic).
Also the field of pervasive computing, focused on integrating computational and communication environments into daily human life, has seen substantial progress. This domain includes a wide spectrum of systems, ranging from high-performance computing (HPC) to resource-constrained IoT devices, with applications spanning smart homes, factories, and hospitals. Systems deployed in this pervasive context must be able to execute complex, high-level tasks. Such execution requires extraction and synthesis of information from diverse structures (tabular, document-based) and modalities (textual, images), coupled with seamless interaction with the surrounding environment. Consequently, LLM-based autonomous agents represent an architectural solution for this necessity. By design, these agents possess the fundamental capabilities to perceive from various information sources, reason the actions to take, and interact with surrounding systems using external actuators (often referred to as tools). This unique combination of capabilities give to agentic systems an enormous potential to operate effectively and autonomously within pervasive computing domains.
However, deploying agents in these environments presents significant computational, and security challenge that necessitate custom solutions. Researchers have proposed various solutions to solve these challenges. This paper analyses strategies for integrating agents into pervasive computing, examines existing applications, identifies persistent challenges, and explores ongoing research efforts to address them. Specifically, the paper is structured as follows:
@@
    First, an introduction to LLM-based agent systems cover their architecture, performance evaluation metrics for accuracy and robustness, and diverse applications.
@@ 2. Agent LLM Architecture
Figure 1. The LLM Agent architecture
In the context of LLM-based agents, a natural language processing pipeline is employed to enable the agent to perceive, reason, and interact with its environment to accomplish specific goals. Perception, the initial stage in this process, involves gathering information about the current state of the environment (wang2024survey). This often occurs through natural language inputs, where users provide instructions or queries (franciscatto2025cbr), or structured text, including HTML or a programming language formats (hu2024automated; zhang2024webpilotversatileautonomousmultiagent). In addition, multimodal LLMs (MLLM), process visual and auditory data, resulting in enhanced environmental comprehension and information use (durante2024agent). Also, to enhance environmental interaction, respect constraints, and achieve goals, LLM-based agents require a robust understanding of implied textual meanings (li2024lasp). Following perception, the agent leverages the Large Language Model’s (LLM) natural language processing capabilities for reasoning and planning. This entails processing the perceived information, interpreting user intent, and analyzing possible courses of action (parmar2025plangen; wang2024survey). A key aspect of this stage is the formulation of strategic plans to achieve desired outcomes. Complex tasks are often decomposed into manageable subtasks to facilitate execution (li2024lasp; parmar2025plangen) such as for robots movements management (yang2025magma), automatic world exploration (wang2023voyager) or code debugging (wang2024openhands). The final stage involves in interacting directly with the environment following the previous reasoning steps (wang2024survey). Actions may include generating natural language responses to the user, providing explanations or completing tasks through dialogue (zhang2024agentic; xi2025rise), invoke external tools or APIs (durante2024agent; schick2023toolformer) or produce executable code (zhang2024aflow). To accomplish all these tasks modern agent architectures typically consist of multiple key modules that work together to create intelligent behavior. A typical LLM-agent architecture has four key modules: a profiling module for defining the agent’s role, a memory module for storing and retrieving past experiences, a planning module for formulating future actions, and an action module for translating decisions into outputs (zhang2024agentic). Also, evaluating this architecture is essential to determine the effectiveness of its modules, especially within the intended application domain. The following sections will explore the LLM architecture, the specific functionalities and interactions of the agent modules, and the evaluation strategies and application domains of these intelligent LLM agents.
@@ 2.1. Large Language Models
Figure 2. The transformer architecture, from (vaswani2017attention)
The core cognitive component, or ”brain”, of intelligent agents are Large Language Models (LLMs). Their fundamental operation rely on ”next token prediction”. This means that given an input sequence of words (tokens), transformed into numerical embeddings, LLMs predict the subsequent token in the sequence, iteratively generating the output. The success of LLMs in this task is largely attributed to their foundation in the Transformer architecture, illustrated in Figure 2. Its design captures long-range dependencies within the text and effectively mimics the complex structures of human language (vaswani2017attention). A typical LLM consists of stacked Transformer layers, each incorporating a multi-head self-attention mechanism and a position-wise fully connected feed-forward network. The multi-head self-attention mechanism allows the model to weigh the importance of different parts of the input sequence when processing each word. It processes the encoded input as a set of key-value pairs (K, V), where both keys and values have a dimension equal to the input sequence length (n). The Transformer employs scaled dot-product attention to calculate how much attention each value should receive based on a query (Q):
@@
LLMs commonly employ this Transformer architecture, scaling it to billions or even trillions of parameters and incorporating various optimization strategies. For instance, decoder-only architectures like GPT-3 and LLaMA predict each next token based on the preceding ones. Conversely, encoder-only architectures such as BERT and RoBERTa prioritise understanding the input text to generate task-specific outputs like labels or token predictions (matarazzo2025survey). In the agentic field, the choice of architecture depends on the required task. Typically, decoder-only architectures are most commonly used when we want an agent to perceive, reason, and interact with the environment. Furthermore, Multimodal Language Models (MLLMs) are increasingly gaining traction for agents interacting in multimodal environments (text, images, video, audio). This is because they allow the application of Transformer architectures across diverse modalities like vision and audio. To enable such multimodal capabilities, MLLMs extend the Transformer architecture by embedding non-text inputs (e.g., image patches or audio tokens) into the same vector space as textual tokens. These embeddings are then processed jointly within shared Transformer layers or through modality-specific experts, such as in Multiway Transformer architectures. Deep modality fusion is achieved via cross-attention or unified encoders, enabling fine-grained interactions between visual and textual features (liang2024survey).
@@ 2.2. Agent Modules
LLM-based autonomous agents are designed to effectively perform diverse tasks by leveraging the capabilities of Large Language Models (LLMs). To achieve this, their architecture typically incorporates 4 key modules (Figure 1): profiling module, memory module, planning module, and action module.
@@ 2.2.1. Profiling
Figure 3. An example of profiling prompt used in (lu2024ai)
The primary purpose of the Profiling Module is to define the role-specific identity of the agent, which significantly influences its behavior and interactions (wang2024survey; zhang2024agentic). This module is crucial for enabling agents to generate more contextually relevant responses from the LLM across different modalities. These profiles are typically added into the model context window. By including role descriptions in the prompt, the profiling module guides the LLM’s reasoning style, response formulation, and interaction strategy (zhang2024agentic). For instance, in (lu2024ai) the agent is profiled to work as phD student that wishes to publish some papers (Figure 3). The profiling module serves as the foundation for agent design, giving a significant influence on the entire agent’s modules ecosystem (wang2024survey). The chosen profile can impact how the agent remembers information (memory module), how it formulates strategies to achieve its goals (planning module) that consequently will impact also the future actions. For example, an agent profiled as ’an expert machine-learning researcher’ (hu2024automated) might prioritise certain experiments over others or selectively memorise results more significant to its research. Similarly, an agent profiled as ’an advanced AI system serving as an impartial judge for intelligent code generation outputs’ (zhuge2024agent) would focus more intently on code details, errors, and bugs than on other capabilities. Agent profiles are commonly implemented by handcrafting specific prompts that describe the desired role and behavior (wang2024survey). However, other strategies can be used, including LLM-generation of profiles and dataset alignment to reflect real-world characteristics (wang2024survey). Combining these strategies can yield additional benefits. Moreover agent profiles can include multiple dimensions, including basic attributes, behavioral patterns, and social information (zhang2024agentic).
@@ 2.2.2. Memory
Large Language Models lack persistent memory across test time interactions, requiring prior exchanges to be explicitly included within the context window (Figure LABEL:fig:memory_schema). This limitation underscores the need for effective memory management strategies in LLM-based agents, particularly for prioritizing relevant information during storage and retrieval. As the agent’s interaction history expands, exceeding the fixed context window’s capacity, summarization or information filtering becomes unavoidable. Consequently, autonomous LLM-based agents rely on a memory module to retain, recall, and reflect upon past observations, thoughts, feedback, and actions. This enables them to better understand the current context, anticipate future needs, correct their action trajectory, adjust their behavior, and maintain consistent and coherent behavior over time (wen2024autodroidv1; xi2025rise; wang2024survey). This module often incorporate different structures to manage information across varying timescales:
@@
  • Short-Term Memory: This type of memory maintains contextually relevant information about recent perceptions, reasoning and actions (the trajectory history) within the LLM context window (zhang2024agentic; shinn2023reflexion). It allows for real-time adaptation during an interactive session. Examples include the internal states maintained by a conversational agent or the intermediate steps during a web research (zhang2024webpilotversatileautonomousmultiagent). However, the limited context window of LLMs poses a challenge for relying solely on short-term memory.
@@
  • Long-Term Memory: In contrast to the real-time encounter observations, which provide short-term memory, the overall encounter history represents long-term memory (zhang2024survey). Long-term memory can provide stable knowledge to complement the flexibility of short-term memory. This module is important because stores both successful experiences, demonstrating correct goal achievement strategies, and unsuccessful experiences, highlighting incorrect paths to avoid. For instance in (wu2024beyond), successful paths are stored as ’Thought Cards’ and retrieved upon encountering a new question with a similar topic or in (zhang2024proagent) the agent store historical events that will be used to predict potential tasks.
@@
  • Hybrid Memory: Many agents use a combination of short-term and long-term memory to leverage the benefits of both (wang2024survey). Short-term memory handles immediate context, while long-term memory provides access to a broader history and consolidated knowledge.
@@
These structures can be stored in various formats, each with its own advantages. One approach is to keep everything in natural language form. This is especially useful for conversational agents, where it’s important to be transparent about the agent’s history. For instance, the method in (shinn2023reflexion) stores past trials as verbal descriptions, helping the model to understand and correct previous attempts. Similarly, in (wang2023voyager) descriptions of skills within the Minecraft game are stored directly as natural language within the agent’s memory. Then, there’s the idea of Embedding Memory. Here, past experiences are transformed into high-dimensional vectors (wang2024survey). This allows for really fast reflection and helps the agent generalise across different situations. It’s especially good for searching through knowledge quickly, based on semantic similarity (zhang2024agentic). Finally databases can be used to store memories. This gives us considerable power to manipulate the data through structured queries. So agents can be used to understand and execute SQL queries in natural language, which lets it interact with the database really effectively (shen2024demonstration).
Three core operations that can be applied to the memory. First writing, which involves persisting relevant information from past interactions into the memory. Key considerations during memory writing include handling duplicated information and managing memory overflow (chen2024llm). Various techniques are used such as summarizing similar information, using a FIFO buffer to overwrite old entries (wang2024survey), truncating historical data that exceeds processing capacity (xi2025rise), and implementing time limits to discard redundant data and prevent excessive memory consumption (zhong2024casit). Next, memory reading is a crucial operation that aims to obtain relevant knowledge from the agent’s memory to adapt its behaviours and inform its next actions (zhang2024agentic). The retrieval of information often considers factors like recency, relevance, and importance of the memories. These factors commonly include the recency of the memory, its relevance to the current situation or query, and its perceived importance (xi2025rise; wang2024survey). Finally, memory reflection allows the agent to analyse, refine, and optimise the accumulated memories (zhang2024agentic). It emulates human cognitive processes of summarizing and inferring more abstract and high-level insights from past experiences. Reflection can occur hierarchically, generating novel based on existing insights (shinn2023reflexion). As mentioned in the previous section, the memory module is not isolated. It is influenced by the agent’s profile, which can determine what types of data are prioritised for storage and retrieval. Furthermore, the memory module plays a critical role in the planning process, providing the necessary context and historical data for the agent to formulate effective future actions. The planning module might retrieve information from memory to inform its reasoning and decision-making.